OpenStudy (anonymous):

Find the roots of the polynomial equation. x2 + 24 = –11x A. 12, 2 B. –8, –3 C. –12, –2 D. 3, 8

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

\[\huge x^2 + 24 = –11x \rightarrow x^2 + 11x+24 =0\] \[\huge \rightarrow x^2 + 8x+3x+24 =0\] \[\huge \rightarrow x(x + 8)+3(x+8) =0\] \[\huge \rightarrow (x + 8)(x+3) =0\] \[\huge \rightarrow (x + 8)=0 or (x+3) =0\] \[\huge \rightarrow x =- 8 or x=-3\] hence x=-8,-3 are the required roots of the given quadratic eq. Thus option B. –8, –3 is the correct answer. @blueweek77

OpenStudy (anonymous):

@blueweek77

OpenStudy (anonymous):

Ok, Thank you so much for explaining it to me!:D